Can you help me understand finding the midpoint at the segment with endpoints (7,-5) and (-3,-9). I beleive that you will add 7 and -3 together and then divide.
Answer by vihits13(20)

ok... to find the x-coordinate of the midpoint, you add both of the x-coordinate given up and then divide that sum by 2.
[7+(-3)]/2=4/2= 2= x coordinate.
you do the same thing to find the y-coordinate of the midpoint: add both of the given y-coordinates up and then divide their sum by 2.
[-5+(-9)]/2=(-14)/2= -7 = y-coordinate
so the midpoint would be at (2,-7)